OpenStudy (usukidoll):

Consider the quadratic equation ax^2+bx+c. Suppose that b and c can only take the values [1,2,3,4,5,6]. For which values of b and c does the equation have real roots?

OpenStudy (usukidoll):

do I have to use the quadratic formula on each combination ?

OpenStudy (usukidoll):

alright..but do I have to plug in the values to the discriminant?

terenzreignz (terenzreignz):

Whether or not it has real roots also depends on the value of a, right? I mean, if a is negative, then it doesn't really matter what b and c are, the quadratic expression WILL have real roots...

OpenStudy (usukidoll):

ummm how did you get b= c/2?

OpenStudy (anonymous):

look at the discriminant b^2-4ac putting the values of abc u can take ur answer

OpenStudy (usukidoll):

ok if we let b=2 and c = 1 b^2-4ac 2^2-4(1)(1) = 4-4 =0

OpenStudy (anonymous):

here we have to consider a too

OpenStudy (usukidoll):

but a is like one on default since b and c is 1-6

terenzreignz (terenzreignz):

Are you absolutely sure that a = 1? Because it doesn't say so... I'm not really a fan of assumptions ^_^

OpenStudy (usukidoll):

so it has to be something...like a b c values that can produce something greater than 0?

OpenStudy (anonymous):

hey poeple a= not equal to one ... for real roots b2-4ac=0 frrom here putting b c we can find a.